# MLVU: A Comprehensive Benchmark for Multi-Task Long Video Understanding

### v2

@@ Abstract
The evaluation of Long Video Understanding (LVU) performance poses an important but challenging research problem. Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances. To address the above problems, we propose a new benchmark, called MLVU [^1] MLVU Repository: https://github.com/JUNJIE99/MLVU. (Multi-task Long Video Understanding Benchmark), for the comprehensive and in-depth evaluation of LVU. MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations. 2) The inclusion of various video genres, e.g., movies, surveillance footage, egocentric videos, cartoons, game videos, etc., which reflects the models’ LVU performances in different scenarios. 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding. The empirical study with 20 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos. Additionally, it suggests that factors such as context length, image-understanding quality, and the choice of LLM backbone can play critical roles in future advancements. We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
@@ 1 Introduction
Large language models (LLMs) are growing into a general solution for numerous AI tasks [6; 40]. In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM. Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1; 39]. Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [51; 27; 9; 21].
@@
However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations. Firstly, the majority of existing video understanding benchmarks are made up of short videos [48; 21; 19; 32; 15], whose lengths can be merely a few seconds. As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities. Secondly, there is a severe lack of diversity for the existing LVU benchmarks in terms of video genres and evaluation tasks. On one hand, the existing benchmarks are usually created based on one type of videos, e.g., movies [36; 22]. On the other hand, it may only contain one specific type of task in each individual benchmark, for example, the captioning task in [48], the temporal perception task in [50], and the action understanding task in [42; 11]. Due to the above limitations, existing benchmarks have difficulty in performing a comprehensive evaluation of LVU performances. Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos. For example, many questions are simply about one single frame in the long videos [36]. Besides, numerous others are about popular movies and celebrities [22], which can be answered directly by MLLMs based on the textual prompts.
@@
Conceptually, the MLLMs are expected to handle any type of long video and accomplish any type of related tasks. Therefore, the evaluation of LVU should emphasize two important properties: length and diversity. Based on such a principle, we propose a novel benchmark called MLVU (Mult-task Long Video Understanding Benchmark), which presents the following critical advantages.
@@
  • It makes a substantial extension for the video length. MLVU is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours. The average video length is about 12 minutes, which makes it much longer than most of the existing benchmarks. Additionally, each video is further segmented so that evaluation tasks can be created w.r.t. different video clips (e.g., summarization for the first 3 minutes, the first 6 minutes, and the entire duration of the video). Therefore, it is able to flexibly evaluate the MLLMs’ performance across different video lengths.
@@
  • It covers a wide variety of video genres. On one hand, it contains different types of real-world videos, such as movies, documentaries, surveillance videos, and ego-centric videos. On the other hand, it also includes typical simulated videos, like games and cartoons. As a result, it is able to comprehensively reflect the MLLMs’ performance across different application scenarios.
  • It introduces diversified evaluation tasks tailored for LVU. There are 9 different tasks in MLVU, which jointly examine a wide range of MLLMs’ key abilities, such as reasoning, captioning, recognition, summarization, etc. Both multi-choice and free-form generation tasks are included in MLVU, which reflect the MLLMs’ performances in handling different forms of tasks. Finally, some of the tasks are designed to leverage global information from entire videos, while others need to utilize proper local information from specific clips. Therefore, it can provide a unified perspective for the degree of completeness and nuance in understanding long videos.
@@
We extensively investigate 20 popular MLLMs with MLVU, which brings in several critical insights. Firstly, long-video understanding remains a technically challenging problem for the existing MLLMs. While GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ achieves the leading performance in the experiment, it only attains an average score of 64.6% in multi-choice tasks. All methods struggle with tasks requiring fine-grained information from entire videos, such as action counting, ordering, and summarization. Additionally, all methods experience significant performance drops as video lengths increase. Secondly, a significant gap persists between open-source and proprietary models. Although some of the latest open-source MLLMs can handle very long videos and excel at single-image tasks, they notably lag behind GPT-4o in understanding long videos. Finally, the empirical results underscore the influential factors in LVU, such as the extension of context length, the improvement of image understanding ability, and the utilization of strong LLM-backbones. In addition to the benchmark’s overall conclusion, the individual tasks enable the fine-grained analysis of MLLMs’ performances in each of the specialized aspects. Therefore, we anticipate the benchmark to assist in improving MLLMs’ long-video understanding capabilities by providing insights into their current strengths and weaknesses.
@@ 2 Related Work
Multimodal Large Langugae Models.
Multimodal large language models (MLLMs) have gained wide interests from both academic and industrial sectors. Recent studies have made significant progresses in this field by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction [25; 56; 7]. Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [53; 29; 20; 49; 23; 21]. However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos. To address this challenge, researchers have explored compact representations of videos. For example, LLaMa-Vid [22] compresses each video frame into two tokens, allowing the model to handle videos several hours long. Methods like MovieChat [36] and MA-LMM [12] introduce specialized memory components for recursive video processing. Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [47; 34; 41]. Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
With the unprecedented interest in MLLMs, the creation of benchmarks for these models has become increasingly emphasized (as advanced by MMMU [51], MME [9], and many other pioneering works). In video understanding, the research community has made significant efforts as well, particularly for short videos. There are specialized benchmarks for temporal perception [50; 43; 28], action understanding [43; 42], video classification [14], video reasoning [46; 45], and video captioning [48; 31]. Recently, MVBench [21] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering. For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks. For example, the LVU dataset [44] presents movie understanding tasks, such as predicting release years and identifying character relationships. Similarly, LLaMA-Vid [22] created a movie question-answering dataset based on MovieNet [13]. Despite using long videos, many questions are about common-sense knowledge of the movies. As a result, they could be directly answered without using the video’s information. In contrast, MovieChat [36] makes use of diversified videos and avoids specific character names or plot details in its questions. Considering that many of the questions target on exact time segments, it potentially degrades the tasks to short-video or image understanding problems. Beyond movies, there are other task-specific benchmarks like EgoSchema [30], which presents video reasoning tasks using first-person footage from Ego4D [11]. However, these specialized benchmarks only focus on one single aspect of MLLMs, rather than offering a comprehensive analysis of long video understanding. Therefore, it remains essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
@@ 3.1 Overview
MLVU is a comprehensive benchmark made up of 2593 evaluation tasks [^2] Only consider the tasks for the entire videos. Others for fine-grained video segments are not accounted. (shown as the right side of Figure 1), which belong to 9 task categories tailored for long-video understanding. These categories include 1) Topic Reasoning, 2) Anomaly Recognition, 3) Video Summarization, 4) Needle Question Answering, 5) Ego Reasoning, 6) Plot Question Answering, 7) Sub-Scene Captioning, 8) Action Count, and 9) Action Order. The benchmark is distinguished by the following features.
Figure 1: Statistical overview of our LVBench dataset. Left: Distribution of Video Duration; Middle: Distribution of Source Types for Long Videos; Right: Quantification of Each Task Type.
Substantial extension of video length.
MLVU is made up of videos of diversified lengths, spanning from 3 min to more than 2 hours (Figure 1 Left). The average length is about 12 min, which is much longer than the existing LVU benchmarks. Besides, each video is further partitioned as incremental segments, e.g., the first 3 min, the first 6 min, and the entire video, where tasks are created for each individual segment. Thus, the MLLMs can be flexibly evaluated across different video lengths.
Diversified video categories.
MLVU offers a comprehensive collection of videos of different categories (Figure 1 Middle). There are typical real-world videos, which includes movies [52; 36], documentaries [36], TV series, egocentric videos [11], and surveillance footage [38]. Meanwhile, there are also important simulated videos from animated series and game videos [8].
Diversified evaluation tasks.
MLVU also provides a diversified collection of evaluation tasks, which are closely related to the common visual capabilities of MLLMs, such as reasoning, captioning, recognition, perception, and summarization (Figure 1 Right). All the tasks are tailored for LVU. That is to say, the tasks need to be solved based on the in-depth understanding of video. Some of tasks are to examine whether the global information from the entire video can be effectively utilized (holistic LVU); while others focus on whether the MLLMs can make precise usage of proper local information within the long video (detail LVU). Additionally, both multi-choice and free-form generation tasks are included in MLVU, which help to examine MLLMs’ capabilities in handling different task formats.
Figure 2: Examples of MLVU. There are nine tasks designed to evaluate the holistic, single-detail, and multi-detail LVU capabilities of MLLMs. The MLLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on the long video input and textual prompt.
@@ 3.2 Construction of MLVU
The evaluation tasks of MLVU which can be categorized into three types: 1) holistic LVU, which needs to be solved by making use of the global information from the entire video; 2) single-detail LVU, which needs to leverage one critical plot within the long video; 3) multi-detail LVU, which calls for the joint utilization of multiple plots within the long video. The construction process of MLVU is discussed w.r.t. the above three categories. To facilitate the discussion, we define ULVC (Universal Long Video Collection) as the universal collection of miscellaneous long videos, which includes movies, documentaries, TV-series, ego-centric videos, surveillance footages, animated series, game videos, et al. (more details about ULVC is presented in the Appendix B).
@@ 3.2.1 Holistic LVU
Topic Reasoning (TR). The topic reasoning task requires MLLMs to respond to questions about the principal subject of a long video, as shown with Figure 2 (a). This includes elements such as the video’s genre, pivotal events, or primary settings. Videos for TR tasks are sourced from movies, documentaries, animated series, egocentric videos, and game videos from ULVC. All questions and answers undergo manual annotation [^3] Detailed information and annotation guidelines for annotators are presented in Appendix E, resulting in a total of 264 questions. TR tasks are formatted as multiple-choice questions, with the model’s performance assessed based on accuracy.
Anomaly Recognition (AR). The anomaly recognition task involves identifying the anomalous behavior within a surveillance footage (Figure 2 b). We leverage the surveillance video clips from UCF Crime dataset [38] for this task. The selected video clips are longer than three minutes. We create 200 questions based on the original annotations provided by the dataset. The AR task is also conducted in the multiple-choice format, whose performance is measured by accuracy.
Video Summarization (VS). This task requires MLLMs to summarize the key events in a long video (Figure 2 c). We select the narrative-rich videos from ULVC for this task, including movies, TV series, documentaries, and animated series. There are 217 selected videos in total, whose summaries are manually annotated. During evaluation, the MLLMs are prompted with "Please summarize the main content of this video". We employ GPT-4 to assess the generated summaries by comparing with the annotation results. Details about annotation and evaluation are presented in Appendix E.3 and F.3.

### v3

# MLVU: Benchmarking Multi-task Long Video Understanding

@@ Abstract
The evaluation of Long Video Understanding (LVU) performance poses an important but challenging research problem. Despite previous efforts, the existing video understanding benchmarks are severely constrained by several issues, especially the insufficient lengths of videos, a lack of diversity in video types and evaluation tasks, and the inappropriateness for evaluating LVU performances. To address the above problems, we propose a new benchmark called MLVU (Multi-task Long Video Understanding Benchmark) for the comprehensive and in-depth evaluation of LVU. MLVU presents the following critical values: 1) The substantial and flexible extension of video lengths, which enables the benchmark to evaluate LVU performance across a wide range of durations. 2) The inclusion of various video genres, e.g., movies, surveillance footage, egocentric videos, cartoons, game videos, etc., which reflects the models’ LVU performances in different scenarios. 3) The development of diversified evaluation tasks, which enables a comprehensive examination of MLLMs’ key abilities in long-video understanding. The empirical study with 23 latest MLLMs reveals significant room for improvement in today’s technique, as all existing methods struggle with most of the evaluation tasks and exhibit severe performance degradation when handling longer videos. Additionally, it suggests that factors such as context length, image-understanding ability, and the choice of LLM backbone can play critical roles in future advancements. We anticipate that MLVU will advance the research of long video understanding by providing a comprehensive and in-depth analysis of MLLMs.
@@ 1 Introduction
Large language models (LLMs) are growing into a general solution for numerous AI tasks [6, 45]. In recent years, it becomes increasingly emphasized to extend LLMs with multi-modal capabilities and thus bring the Multi-modal LLM, namely, MLLM. Remarkably, it has been made possible for today’s MLLMs to perceive information in texts, images, videos, etc., and solve complicated problems in physical environments [1, 44]. Along with the development of MLLMs, new benchmarks are continuously created to facilitate comprehensive and in-depth analysis of MLLMs [57, 32, 12, 26].
@@
However, it remains a great challenge to evaluate the MLLMs’ long-video understanding (LVU) performances given the following limitations. Firstly, the majority of existing video understanding benchmarks are made up of short videos [52, 26, 22, 36, 19], whose lengths can be merely a few seconds. As a result, they are insufficient to reflect the MLLMs’ long-video understanding capabilities. Secondly, there is a notable lack of diversity in both video genres and evaluation tasks. Existing benchmarks often concentrate on a single video type, such as egocentric videos [15, 34], or focus on one specific task, like captioning [52]. These limitations hinder comprehensive evaluation of LVU capabilities. Last but not least, many previous evaluation tasks are not properly designed for LVU, as they can be solved without using the complex information from long videos. For example, many questions are simply about one single frame in the long videos [41, 60]. Besides, numerous others are about popular movies and celebrities [27, 13], which can be answered directly by MLLMs based on the textual prompts.
@@
Conceptually, MLLMs are expected to handle any type of long video and accomplish any related tasks. Therefore, the evaluation of LVU should emphasize two important properties: length and diversity. Furthermore, it is crucial that the evaluation tasks are specifically designed to leverage the complex information inherent in long videos, addressing the shortcomings of previous benchmarks. Based on these principles, we propose a novel benchmark called MLVU (Mult-task Long Video Understanding Benchmark), which presents the following critical advantages.
Benchmarks #Videos #QA Pairs Len. (s) Close-Ended Open-Ended Various Genres Multi-Level Multi-Dimension Referring QA NExT-QA [50] 1,000 8,564 39.5 TVQA [21] 15,253 15,253 11.2 MSRVTT-QA [52] 2,900 72,821 15.2 MVBench [26] 3,641 4,000 16.0 Movie101 [58] 101 - 6144 EgoSchema [34] 5,063 5,063 180 MovieChat-1K [41] 130 1,950 500 Video-MME [13] 900 2,700 1024 LongVideoBench [49] 3,763 6,678 473 MLVU 1,730 3,102 930
Table 1: Comparison of MLVU with existing benchmarks, including the number of videos (#Videos), number of QA pairs (#QA pairs), average video length (Len.), presence of Close-Ended tasks, presence of Open-Ended tasks, inclusion of various video genres (Various Genres), coverage of multiple duration levels (Multi-Level), inclusion of multiple dimensions of LVU tasks (Multi-Dimension), and questions involving local information with clear referring context rather than direct timestamps [41] or well-known narrative elements [17, 27] (Referring QA). The first block represents short video understanding benchmarks, and the second block represents long video understanding benchmarks. denotes work concurrent with MLVU.
@@
  • It makes a substantial extension for the video length. MLVU is created based on long videos of diversified lengths, ranging from 3 minutes to 2 hours. The average video length is about 15 minutes, which makes it much longer than most of the existing benchmarks. Additionally, each video is further segmented so that evaluation tasks can be created w.r.t. different video clips (e.g., summarization for the first 3 minutes, the first 6 minutes, and the entire duration of the video). Therefore, it is able to flexibly evaluate the MLLMs’ performance across different video lengths.
@@
  • It encompasses a wide variety of video genres. MLVU includes diverse real-world videos, such as movies, life records, and egocentric videos. Additionally, it features typical simulated videos like games and cartoons. This diversity allows for a comprehensive assessment of MLLMs’ performance across various application scenarios.
  • It introduces diversified evaluation tasks tailored for LVU. MLVU comprises 9 distinct tasks that collectively assess a wide range of MLLMs’ LVU capabilities. On one hand, it includes both multiple-choice and open-ended generation tasks, reflecting the models’ performance in handling different task formats. On the other hand, some tasks are designed to leverage global information from entire videos, while others require the use of specific local information from certain clips. Moreover, all questions involving local information are annotated with unambiguous context, requiring MLLMs to accurately locate or infer the appropriate clips within long videos.
@@
Table 1 shows that MLVU provides a more comprehensive evaluation of LVU compared to existing and concurrent benchmarks. We extensively investigate 23 popular MLLMs with MLVU, which brings in several critical insights. Firstly, long-video understanding remains a technically challenging problem for the existing MLLMs. While GPT-4o [^1] https://openai.com/index/hello-gpt-4o/ achieves the leading performance in the experiment, it only attains an average score of 54.5% in multi-choice tasks. All methods struggle with tasks requiring fine-grained information from entire videos, such as action counting, ordering, and summarization. Secondly, recent open-source long video MLLMs have made significant strides in LVU [60, 11, 40]. These advancements have improved the models’ capability to process extended visual sequences, thereby closing the gap with leading proprietary models in recent months. Finally, the empirical results underscore influential factors in LVU, such as the extension of context length, the improvement of image understanding ability, and the utilization of strong LLM-backbones. In addition to the benchmark’s overall conclusion, individual tasks enable fine-grained analysis of MLLMs’ performances in each specialized aspects. Therefore, we anticipate the benchmark to assist in improving MLLMs’ long-video understanding capabilities by providing insights into their current strengths and weaknesses.
@@ 2 Related Work
Figure 1: Statistical Overview of our MLVU benchmark. Left: Video genres included in MLVU; Top Right: Distribution of video duration; Bottom Right: Task types and their counts in MLVU.
Multimodal Large Language Models.
Multimodal large language models (MLLMs) have attracted significant interest from both academia and industry. Recent advancements in this field have been achieved by integrating LLM backbones with visual encoders and adapters, and fine-tuning the entire architecture through visual instruction tuning [29, 63, 8]. Based on the same philosophy, MLLMs have been further developed for video processing using video instruction datasets and specialized video adapters [59, 33, 25, 54, 28, 26]. However, most existing models are optimized for short videos, typically under one minute, due to the difficulty in establishing sufficient context for longer videos. To address this challenge, researchers have explored compact video representations or extended the context length of MLLMs. For instance, LLaMa-Vid [27] compresses each video frame into two tokens, enabling the model to handle videos several hours long. Methods like MovieChat [41] and MA-LMM [16] introduce specialized memory components for recursive video processing. Furthermore, approaches such as LWM [30], LongVA [60], and Video-XL [40] are designed to extend the context length of MLLMs, facilitating the processing of longer video inputs. Additionally, it is also explored to make selective usage of frames or clips from long videos based on retrievers or agents [53, 38, 46]. Despite these progresses, it remains an open problem for MLLMs to effectively handle long videos.
@@
With the unprecedented interest in MLLMs, the creation of benchmarks for these models has become increasingly emphasized (as advanced by MMMU [57], MME [12], and many other pioneering works). In video understanding, the research community has made significant efforts as well, particularly for short videos. There are specialized benchmarks for temporal perception [56, 48], action understanding [48, 47], video classification [18], video reasoning [51, 50], and video captioning [52, 35]. Recently, MVBench [26] provides a comprehensive short-video benchmark to evaluate general capabilities via question-answering. For long video understanding, people seek to leverage long-form videos, like movies, to create benchmarks. For example, LLaMA-Vid [27] developed a movie question-answering dataset based on MovieNet [17]. Despite using long videos, many questions focus on well-known narrative elements, allowing them to be answered without analyzing the video’s content. In contrast, MovieChat [41] avoids specific character names or plot details in its questions. However, since each question provides a specific timestamp, the tasks can be reduced to short-video or image understanding problems. Beyond movies, there are task-specific benchmarks like EgoSchema [34], which presents video reasoning tasks using first-person footage from Ego4D [15]. These specialized benchmarks, however, focus on a single aspect of MLLMs rather than offering a comprehensive analysis of long video understanding. Therefore, it is essential to develop a comprehensive benchmark with carefully designed tasks to effectively evaluate MLLMs’ capabilities in understanding long videos.
@@ 3.1 Overview
MLVU is a multi-task benchmark consisting of 3,102 questions across 9 categories, specifically designed for long video understanding. It is divided into a dev set and a test set, containing 2,593 and 509 questions, respectively. The benchmark is distinguished by the following features.
Diversified Video Categories. MLVU offers a comprehensive collection of videos across various categories (Figure 1 Left). These include typical real-world videos such as movies, documentaries, TV series, egocentric videos, life records, sports, tutorials, and surveillance footage. Additionally, it features significant simulated videos from animated series and game videos.
Substantial Extension of Video Length. MLVU is made up of videos of diversified lengths, spanning from 3 min to more than 2 hours (Figure 1 Top Right). Besides, each video is further partitioned as incremental segments, e.g., the first 3 min, the first 6 min, and the entire video, where tasks are created for each individual segment. Thus, the MLLMs can be flexibly evaluated across different video lengths.
Diversified Evaluation Tasks. MLVU also provides a diverse array of evaluation tasks, which are closely aligned with the common visual capabilities of MLLMs, such as reasoning, captioning, recognition, perception, and summarization (Figure 1 Bottom Right). All the tasks are tailored for LVU. That is to say, the tasks need to be solved based on the in-depth understanding of video. Some of tasks are to examine whether the global information from the entire video can be effectively utilized (holistic LVU); while others focus on whether the MLLMs can make precise usage of proper local information within the long video (detail LVU). Additionally, both multi-choice and free-form generation tasks are included in MLVU, which help to examine MLLMs’ capabilities in handling different task formats.
@@ 3.2 Construction of MLVU
The evaluation tasks of MLVU can be categorized into three types: 1) holistic LVU, which needs to be solved by making use of the global information from the entire video; 2) single-detail LVU, which relies on leveraging one critical plot within the long video; and 3) multi-detail LVU, which necessitates the joint utilization of multiple plots within the long video. The construction process of MLVU is discussed w.r.t the above three categories. To facilitate the discussion, we define ULVC (Universal Long Video Collection) as the universal collection of long videos from various sources (more details about ULVC are presented in Appendix C).
@@ 3.2.1 Holistic LVU
Topic Reasoning (TR). The topic reasoning task requires MLLMs to respond to questions about the principal subject of a long video, as shown with Figure 2 (a). This includes elements such as the video’s genre, pivotal events, or primary settings. All questions and answers undergo manual annotation [^2] Detailed information and annotation guidelines for annotators are presented in Appendix F., resulting in a total of 355 questions. TR tasks are formatted as multiple-choice questions, with the model’s performance assessed based on accuracy.
Anomaly Recognition (AR). The anomaly recognition task involves identifying the anomalous behavior within a surveillance footage (Figure 2 b). We leverage the surveillance video clips from UCF Crime dataset [43] for this task. The selected video clips are longer than three minutes. We create 239 questions based on the original annotations provided by the dataset. The AR task is also conducted in the multiple-choice format, whose performance is measured by accuracy.
Video Summarization (VS). This task requires MLLMs to summarize the key events in a long video (Figure 2 c). We select the narrative-rich videos from ULVC for this task, including movies, TV series, documentaries, life records, and animated series. There are 257 selected videos in total, whose summaries are manually annotated. During evaluation, the MLLMs are prompted with "Please summarize the main content of this video". We employ GPT-4 to assess the generated summaries by comparing with the annotation results. Details about annotation and evaluation are presented in Appendix F.3 and G.3.
Figure 2: Examples of MLVU. There are nine tasks designed to evaluate the holistic, single-detail, and multi-detail LVU capabilities of MLLMs. The MLLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on the long video input and textual prompt. For multiple-choice questions, we set 4 candidates in the dev set and 6 candidates in the test set.